Ammonia gas dissolved in 1.0 x 10^-4 M solution of copper sulfate to give an equilibrium concentration of ammonia 1.60 x 10^-3 M.

Calculate the concentration of copper ions in the solution. Kf=5.0 x 10^13
Answer #1

Answer: The reaction of NH3 and Cu2+ ions from CuSO4 takes place as follows:

Cu2+(aq) + 4NH3(aq) ⇌ [Cu(NH3)4]2+(aq)

Given the equilibrium constant for above reaction    Kf = 5.0×1013

The calculations are made in the following table:

         Cu2+(aq)                +                  4NH3(aq)         ⇌           [Cu(NH3)4]2+(aq)

Initial conc. (M)

1.0×10-4

1.60×10-3 + 4x

0

Change in conc. (M)

- x

- 4x

+ x

Equilibrium conc. (M)

1.0×10-4- x

1.60×10-3

+ x

K [Cu(NH3)22-) [Cu]^[NH] 5x 1013 (1x 104 - x) *(1.60 x 10-3

(1×10-4x) (5 × 1013) × (1.60×10-3)4 = x

327.68 (1×10-4x) = x

Or,     x + 327.68 x = 327.68×10-4

Or,     328.68 x = 327.68×10-4

Or,   x = (327.68×10-4) / (328.68)     

x = 0.00009

The concentration for Cu2+ in solution is (1.0×10-4 - 0.00009) = 0.00001 M ≈ 1.0×10-5